A random sample of 90 students at a private high school is tested and a mean SAT score of 1030 is obtained. The average score for public high school student is 1000 (σ= 200).

Compute Z obt
1
Expert's answer
2011-05-24T10:40:39-0400
Zobt = (x - μ)/(σ /√ n) = (1030 - 1000)/(200/√90) = 1.42